8. A speeding driver exits a highway onto an exit ramp going at 50 m/s. The exit ramp is in the shape of a full circle with a diameter of 100 m. Keeping his speed constant, how long does it take him to drive the full length of the exit ramp?

9. For the driver in problem 8, if his mass is 100 kg, and he is pressed sideways against the driver’s side door while he is on the exit ramp, how much normal force must the door exert to keep him inside the car
Answer #1

8) given the shape is a full circle and and the diameter is 100m then the circumferrence is = 314m

then the time taken= D/V = 314/ 50 = 6.28s

9) the force acting on the driver when he is pressed to the side door is same as centrifugal of his weight then in this case the = V/R = 50/50 = 1 then the centrifugal force formula is =100Kg *50m *12 = 5000N acts on the driver when he is pressed against the door